# Runyang You1    Yongqi Li1   Xinyu Lin2   Xin Zhang1 Wenjie Wang3   Wenjie Li1   Liqiang Nie4 1

### v1

@@ 5.3 Ablation Study
w/o Reasoning 0.017 5670.0175670.017\,5670.017 567 0.012 0610.0120610.012\,0610.012 061 0.029 6030.0296030.029\,6030.029 603 0.015 321 020 0110.0153210200110.015\,321\,020\,0110.015 321 020 011 0.051 0740.0510740.051\,0740.051 074 0.020 0110.0200110.020\,0110.020 011 0.046 8750.0468750.046\,8750.046 875 0.032 141 739 768 641 3340.0321417397686413340.032\,141\,739\,768\,641\,3340.032 141 739 768 641 334 0.069 196 428 571 428 580.069196428571428580.069\,196\,428\,571\,428\,580.069 196 428 571 428 58 0.039 296 044 879 371 210.039296044879371210.039\,296\,044\,879\,371\,210.039 296 044 879 371 21 0.094 494 047 619 047 620.094494047619047620.094\,494\,047\,619\,047\,620.094 494 047 619 047 62 0.045 616 438 834 085 350.045616438834085350.045\,616\,438\,834\,085\,350.045 616 438 834 085 35 0.027 712 171 311 604 4720.0277121713116044720.027\,712\,171\,311\,604\,4720.027 712 171 311 604 472 0.017 357 424 748 561 4170.0173574247485614170.017\,357\,424\,748\,561\,4170.017 357 424 748 561 417 0.044 087 545 268 461 660.044087545268461660.044\,087\,545\,268\,461\,660.044 087 545 268 461 66 0.022 660 610 718 420 0330.0226606107184200330.022\,660\,610\,718\,420\,0330.022 660 610 718 420 033 0.074 791 371 437 568 90.07479137143756890.074\,791\,371\,437\,568\,90.074 791 371 437 568 9 0.030 337 079 249 673 180.030337079249673180.030\,337\,079\,249\,673\,180.030 337 079 249 673 18
w/o $R_{d}$ 0.019 8440.0198440.019\,8440.019 844 0.012 4110.0124110.012\,4110.012 411 0.033 8320.0338320.033\,8320.033 832 0.016 4140.0164140.016\,4140.016 414 0.055 9530.0559530.055\,9530.055 953 0.022 4250.0224250.022\,4250.022 425 0.052 0830.0520830.052\,0830.052 083 0.033 7820.0337820.033\,7820.033 782 0.076 6370.0766370.076\,6370.076 637 0.040 4270.0404270.040\,4270.040 427 0.097 4320.0974320.097\,4320.097 432 0.048 6430.0486430.048\,6430.048 643 0.030 2310.0302310.030\,2310.030 231 0.019 5720.0195720.019\,5720.019 572 0.048 6540.0486540.048\,6540.048 654 0.025 360.025360.025\,360.025 36 0.079 830.079830.079\,830.079 83 0.033 2230.0332230.033\,2230.033 223
w/o $R_{c}$ 0.024 3980.0243980.024\,3980.024 398 0.016 0360.0160360.016\,0360.016 036 0.039 3620.0393620.039\,3620.039 362 0.020 8190.0208190.020\,8190.020 819 0.060 5070.0605070.060\,5070.060 507 0.025 8470.0258470.025\,8470.025 847 0.054 3150.0543150.054\,3150.054 315 0.038 2270.0382270.038\,2270.038 227 0.077 3810.0773810.077\,3810.077 381 0.045 5870.0455870.045\,5870.045 587 0.101 190.101190.101\,190.101 19 0.051 4650.0514650.051\,4650.051 465 0.031 6490.0316490.031\,6490.031 649 0.020 2010.0202010.020\,2010.020 201 0.053 3770.0533770.053\,3770.053 377 0.026 3580.0263580.026\,3580.026 358 $0.0814$ 0.035 5220.0355220.035\,5220.035 522
$\text{R}^{2}\text{ec}$ 0.026 350.026350.026\,350.026 35 0.016 0810.0160810.016\,0810.016 081 0.039 6880.0396880.039\,6880.039 688 0.020 3050.0203050.020\,3050.020 305 0.061 4830.0614830.061\,4830.061 483 0.025 7280.0257280.025\,7280.025 728 0.058 780.058780.058\,780.058 78 0.038 8350.0388350.038\,8350.038 835 0.080 3570.0803570.080\,3570.080 357 0.045 6550.0456550.045\,6550.045 655 0.108 630.108630.108\,630.108 63 0.052 5010.0525010.052\,5010.052 501 0.032 593 292 394 898 440.032593292394898440.032\,593\,292\,394\,898\,440.032 593 292 394 898 44 0.020 513 333 770 876 2880.0205133337708762880.020\,513\,333\,770\,876\,2880.020 513 333 770 876 288 0.053 062 509 840 969 9260.0530625098409699260.053\,062\,509\,840\,969\,9260.053 062 509 840 969 926 0.027 110.027110.027\,110.027 11 0.085 3410.0853410.085\,3410.085 341 0.036 2930.0362930.036\,2930.036 293
@@
We conducted ablation studies by evaluating the following variants: 1) “w/o Reasoning”. We eliminated reasoning tokens from prompts and trained exclusively with in-batch contrastive loss; 2) “w/o $R_{c}$”. We keep only the discrete ranking reward $R_{d}$ and removed the continuous similarity reward $R_{c}$ in Eqn.(5); and 3) “w/o $R_{d}$”. We removed the discrete ranking reward $R_{d}$ from Eqn.(5). The results are summarized in Table 2, and several observations stand out.
@@
1) It is found that $\text{R}^{2}\text{ec}$ achieves an average improvement of roughly 15% across all metrics compared to w/o Reasoning. These substantial gains demonstrate that our designed optimization have enabled $\text{R}^{2}\text{ec}$ to better leverage test-time scaling to deliver significantly stronger recommendation performance. 2) As revealed that w/o $R_{c}$ (using only $R_{d}$) consistently outperforms w/o $R_{d}$ (using only $R_{c}$), this indicates that adopting reward signal that directly reflects evaluation result is crucial for training, while the continuous reward $R_{c}$, despite offering finer granularity, fails to provide meaningful distinctions and instead introduces noise that leads to suboptimal performance. 3) By fusing $R_{d}$ with a small weight on $R_{c}$, our approach preserves the task alignment of ranking rewards while benefiting from the supplementary signal of the continuous term. As a result, $\text{R}^{2}\text{ec}$ achieves optimal performance on nearly all metrics.
5.4 In-depth Analysis
@@ 5.4.2 Analysis on Trajectory Sampling
For $\text{R}^{2}\text{ec}$’s training, we performed trajectory sampling via the temperature $\tau$ and top-$K$ sampling, which influence the stochasticity and diversity of the generated samples. To quantify their impact, we varied $\tau$ and $K$, and the results are presented in Figure 3. 1) It is observed that increasing the temperature produces longer resoning and boosts recommendation performance, i.e., , NDCG@5. A higher temperature introduces greater sampling entropy, allowing the model to explore a wider range of reasoning trajectories. 2) Conversely, it is found that increasing top-$K$ actually shortens the reasoning length and generally leads to a decline in recommendation performance. This is because a larger top-$K$ enlarges the candidate token set, which counterintuitively mitigates length hacking yet excessively large $K$ reintroduces noisy and low-quality samples.
@@ 5.4.3 Analysis on Group Size
During training, we sampled a group of trajectories to estimate their advantages. Therefore, we conducted experiments to analyze how varying group sizes impact performance. The results are outlined in Figure 3, and we gained several key findings as follows: 1) It is observed that performance improves for both backbones as group size increases, but the rate of improvement gradually slows down. While a larger group size generally leads to more explored paths, it also raises the training cost. These results suggest that selecting a group size of 6 or 8 is sufficient, and further increasing the group size is unnecessary. 2) Comparing the two backbones, we found a difference in sensitivity to group size. Qwen’s performance at a group size of 1 lags significantly behind Gemma’s, but it improves rapidly as the group size increases. Gemma performs well even with smaller groups, likely because its pretraining exposed it to a broader range of reasoning scenarios, thereby equipping it with stronger initial reasoning-for-recommendation capabilities.
5.4.4 More Analysis
Due to the scope limitation, we detail more analysis on the case study in Appendix H, and the analysis on the efficiency in Appendix G, respectively. We also discuss the limitations in Appendix F.
@@ 6 Conclusion and Future Work
[…]
Appendix A Dataset
@@
Following the temporal-truncation protocol introduced by [3, 2], we construct three subsets—CDs and Vinyl (CDs), Video Games (Games), and Musical Instruments (Instruments) – from the lastest public Amazon review dataset [^1] https://amazon-reviews-2023.github.io/index.html. spanning from May 1996 to October 2023. For each domain, we begin with the most recent year of interactions (October 2022 – October 2023) and, if the resulting number of valid items is insufficient, iteratively roll the time window backward month-by-month until 10k items are obtained. We omit the 5-core filter so as to retain the nature behaviour characteristic of recommendation scenarios. Each user’s interaction history is then chronologically sorted and truncated to the latest 20 actions, yielding fixed-length sequences for all subsequent modelling and evaluation. Finally the dataset is split with 80% training 10% validation 10% test. The resulting statistics are listed in Table 3.
Appendix B Baselines

### v3

@@ 4.3 Ablation Study
w/ ClsHead $0.0044$ $0.0023$ $0.0102$ $0.0033$ $0.0179$ $0.0067$ $0.0030$ $0.0025$ $0.0045$ $0.0027$ $0.0095$ $0.0044$ $0.0012$ $0.0008$ $0.0022$ $0.0011$ $0.0133$ $0.0032$
w/o Reasoning 0.017 5670.017\,567 0.012 0610.012\,061 0.029 6030.029\,603 0.015 321 020 0110.015\,321\,020\,011 0.051 0740.051\,074 0.020 0110.020\,011 0.046 8750.046\,875 0.032 141 739 768 641 3340.032\,141\,739\,768\,641\,334 0.069 196 428 571 428 580.069\,196\,428\,571\,428\,58 0.039 296 044 879 371 210.039\,296\,044\,879\,371\,21 0.094 494 047 619 047 620.094\,494\,047\,619\,047\,62 0.045 616 438 834 085 350.045\,616\,438\,834\,085\,35 0.027 712 171 311 604 4720.027\,712\,171\,311\,604\,472 0.017 357 424 748 561 4170.017\,357\,424\,748\,561\,417 0.044 087 545 268 461 660.044\,087\,545\,268\,461\,66 0.022 660 610 718 420 0330.022\,660\,610\,718\,420\,033 0.074 791 371 437 568 90.074\,791\,371\,437\,568\,9 0.030 337 079 249 673 180.030\,337\,079\,249\,673\,18
w/o $R_{d}$ 0.019 8440.019\,844 0.012 4110.012\,411 0.033 8320.033\,832 0.016 4140.016\,414 0.055 9530.055\,953 0.022 4250.022\,425 0.052 0830.052\,083 0.033 7820.033\,782 0.076 6370.076\,637 0.040 4270.040\,427 0.097 4320.097\,432 0.048 6430.048\,643 0.030 2310.030\,231 0.019 5720.019\,572 0.048 6540.048\,654 0.025 360.025\,36 0.079 830.079\,83 0.033 2230.033\,223
w/o $R_{c}$ 0.024 3980.024\,398 0.016 0360.016\,036 0.039 3620.039\,362 0.020 8190.020\,819 0.060 5070.060\,507 0.025 8470.025\,847 0.054 3150.054\,315 0.038 2270.038\,227 0.077 3810.077\,381 0.045 5870.045\,587 0.101 190.101\,19 0.051 4650.051\,465 0.031 6490.031\,649 0.020 2010.020\,201 0.053 3770.053\,377 0.026 3580.026\,358 $0.0814$ 0.035 5220.035\,522
$\text{R}^{2}\text{ec}$ 0.026 350.026\,35 0.016 0810.016\,081 0.039 6880.039\,688 0.020 3050.020\,305 0.061 4830.061\,483 0.025 7280.025\,728 0.058 780.058\,78 0.038 8350.038\,835 0.080 3570.080\,357 0.045 6550.045\,655 0.108 630.108\,63 0.052 5010.052\,501 0.032 593 292 394 898 440.032\,593\,292\,394\,898\,44 0.020 513 333 770 876 2880.020\,513\,333\,770\,876\,288 0.053 062 509 840 969 9260.053\,062\,509\,840\,969\,926 0.027 110.027\,11 0.085 3410.085\,341 0.036 2930.036\,293
@@
We conducted ablation studies by evaluating the following variants: 1) “w/ ClsHead”, an alternative design that uses classification head to the hidden state of reasoning tokens, instead of using the unified reasoning–recommendation formulation. 2) “w/o Reasoning”, where reasoning tokens are removed from prompts and the model is trained solely with in-batch contrastive loss; 3) “w/o $R_{c}$”, which retains only the discrete ranking reward $R_{d}$ while removing the continuous similarity reward $R_{c}$ in Eqn. (5); 4) “w/o $R_{d}$”, which removes the discrete ranking reward $R_{d}$ from Eqn. (5); and The results are summarized in Table 2, and several observations stand out.
@@
1) It is found that $\text{R}^{2}\text{ec}$ achieves an average improvement of roughly 15% across all metrics compared to w/o Reasoning. These substantial gains demonstrate that our designed optimization have enabled $\text{R}^{2}\text{ec}$ to better leverage test-time scaling to deliver significantly stronger recommendation performance. 2) The w/ ClsHeadvariant substantially underperforms our tightly-coupled architecture, highlighting the necessity of reasoning–recommendation co-optimization. 3) As revealed that w/o $R_{c}$ (using only $R_{d}$) consistently outperforms w/o $R_{d}$ (using only $R_{c}$), this indicates that adopting reward signal that directly reflects evaluation result is crucial for training, while the continuous reward $R_{c}$, despite offering finer granularity, fails to provide meaningful distinctions and instead introduces noise that leads to suboptimal performance. 4) By fusing $R_{d}$ with a small weight on $R_{c}$, our approach preserves the task alignment of ranking rewards while benefiting from the supplementary signal of the continuous term. As a result, $\text{R}^{2}\text{ec}$ achieves optimal performance on nearly all metrics.
4.4 Analysis on Optimization
This subsection presents analyses of key optimization strategies in RecPO, evaluating their impact on final performance.
@@ 4.4.2 Analysis on Trajectory Sampling
During RecPO, we performed trajectory sampling via the temperature $\tau$ and top-$K$ sampling, which influence the stochasticity and diversity of the generated samples. To quantify their impact, we varied $\tau$ and $K$, and the results are presented in Figure 3. 1) It is observed that increasing the temperature produces longer resoning and boosts recommendation performance, i.e., , NDCG@5. A higher temperature introduces greater sampling entropy, allowing the model to explore a wider range of reasoning trajectories. 2) Conversely, it is found that increasing top-$K$ actually shortens the reasoning length and generally leads to a decline in recommendation performance. This is because a larger top-$K$ enlarges the candidate token set, which counterintuitively mitigates length hacking yet excessively large $K$ reintroduces noisy and low-quality samples.
@@ 4.4.3 Analysis on Group Size
During RecPO, we sampled a group of trajectories to estimate their advantages. Therefore, we conducted experiments to analyze how varying group sizes impact performance. The results are outlined in Figure 3, and we gained several key findings as follows: 1) It is observed that performance improves for both backbones as group size increases, but the rate of improvement gradually slows down. While a larger group size generally leads to more explored paths, it also raises the training cost. These results suggest that selecting a group size of 6 or 8 is sufficient, and further increasing the group size is unnecessary. 2) Comparing the two backbones, we found a difference in sensitivity to group size. Qwen’s performance at a group size of 1 lags significantly behind Gemma’s, but it improves rapidly as the group size increases. Gemma performs well even with smaller groups, likely because its pretraining exposed it to a broader range of reasoning scenarios, thereby equipping it with stronger initial reasoning-for-recommendation capabilities.
4.4.4 Analysis on Embedding Strategy
We further examined different strategies for obtaining sequence-level embeddings. Specifically, we compared the last hidden state used in R2ec against max pooling, mean pooling, and adding a special-token under identical settings. As reported in Table 3, the last hidden state yields the best overall performance. Special token strategies yield weaker performance in our setting, likely because we adopt LoRA fine-tuning and no large-scale optimization on the LLLM backbone; it is thus challenging for the model to effectively learn to generate and utilize newly introduced tokens for global sequence representation.
Table 3: Comparison of sequence embedding extraction strategies.
Method N@5 N@10 H@5 H@10
Special token 0.0353 0.0432 0.0509 0.0737
Mean pooling 0.0368 0.0453 0.0528 0.0796
Max pooling 0.0382 0.0445 0.0595 0.0729
Last hidden state (ours) 0.0388 0.0457 0.0588 0.0804
4.5 Analysis on Reasoning Behavior : How Does $\text{R}^{2}\text{ec}$ Reason to Recommend?
How exactly does $\text{R}^{2}\text{ec}$ transform raw behavioral descriptions into insightful recommendations? While the previous sections establish that reasoning substantially enhances recommendation accuracy, the underlying decision process—how the model internally reasons to reach its conclusions—remains less understood. To demystify this process, we conduct a systematic qualitative and quantitative analysis to uncover the reasoning strategies $\text{R}^{2}\text{ec}$ employs across different domains and user contexts.
Qualitative Analysis.
We first engaged in extensive discussions with colleagues to collaboratively analyze and summarize the diverse reasoning behaviors exhibited by $\text{R}^{2}\text{ec}$. Through this process, we identified seven distinct reasoning strategies of $\text{R}^{2}\text{ec}$ for recommendation:
  1. 1.
    Attribute Abstraction: Identifying underlying attributes of purchased items.
  2. 2.
    Pattern Recognition & Clustering: Detecting frequent item patterns to form user-interest clusters.
  3. 3.
    Scenario/Role-Based Reasoning: Inferring user roles or specific use cases.
  4. 4.
    Temporal Reasoning: Leveraging purchase timing and order for contextual understanding.
  5. 5.
    Self-Explanation: Providing explicit rationales for recommendations.
  6. 6.
    Negative Preference Exclusion: Avoiding previously negatively rated item types.
  7. 7.
    Multi-Objective Balancing: Managing trade-offs between conflicting user objectives.
These reasoning strategies illustrate how reasoning aids recommendation. Equipped with reasoning, large recommender models can distill salient user preferences from noisy or complex behavioral data, abstract beyond surface-level co-occurrence, and produce more contextually appropriate and interpretable recommendations.
Figure 4: Distribution of reasoning behaviors across datasets. Each bar represents the proportion of reasoning outputs exhibiting a given reasoning behavior within a dataset.
Quantitative Analysis.
Building on the above findings, we then randomly sampled 100 reasoning outputs from each test set. These samples were annotated and analyzed through a combination of detailed human evaluation and automated assessment with GPT-4.1, ensuring both depth and consistency. Figure 4 visualizes the proportion of sampled reasoning outputs within each dataset that exhibit each reasoning behavior, which reveals two key trends. (1) The distribution of reasoning behaviors varies across datasets, indicating that $\text{R}^{2}\text{ec}$ self-adapts its reasoning patterns to domain characteristics and user–item interactions. (2) Even within a single domain, $\text{R}^{2}\text{ec}$ flexibly selects different reasoning strategies depending on the specific user sequence. Overall, these results demonstrate that $\text{R}^{2}\text{ec}$ does not follow a fixed reasoning template but instead employs context-aware reasoning tailored to both dataset and user, yielding improved recommendation performance and interpretability.
4.6 Analysis on Inference Efficiency
{wrapfloat}
tabler0.3 Average inference latency (in seconds) across models.
Method Latency (s)
SASRec 0.014
LangPTune 1.90
D3 4.62
LLaRA 5.23
[…]
To quantify the efficiency gains of our unified model $\text{R}^{2}\text{ec}$ over both traditional and LLM-based recommenders, we conducted comparative latency tests. All inferences were performed on a single NVIDIA RTX 3090 GPU using identical input prompts, maximum token lengths, and candidate sets. We executed 100 queries per run at a batch size of 1 and averaged the results over 3 independent runs.
As shown in Table 4.6, 1) $\text{R}^{2}\text{ec}$ achieves competitive inference efficiency among LLM-based recommenders, outperforming comparable methods such as LangPTune, D3, and LLaRA. Such benefits stem from the use of scalable item embeddings, which avoid expensive autoregressive decoding over large item vocabularies and the design of an unified architecture. 2) When deployed with the VLLM inference framework, $\text{R}^{2}\text{ec}$ significantly narrowed the efficiency gap with traditional sequential models like SASRec. These results demonstrate that $\text{R}^{2}\text{ec}$ successfully balances expressive power and computational efficiency.
4.7 More Analysis
Due to space constraints, further extended analyses on Scaling to Larger Backbones and Cross-Domain Robustness are reported in Appendix I, while Appendix J provides representative case studies. We also discuss the limitations of our approach in Appendix H.
@@ 5 Conclusion and Future Work
In this study, we investigate the integration of reasoning into large recommender models by introducing $\text{R}^{2}\text{ec}$, a unified large recommender model with intrinsic reasoning capabilities, trained with RecPO, without reliance on human-annotated reasoning annotations. Extensive experiments and analysis demonstrated that $\text{R}^{2}\text{ec}$achieves both higher recommendation quality and lower inference latency compared to existing reasoning-augmented and conventional LLM-based recommenders. Taken together, these findings highlight the importance of tightly coupling reasoning and recommendation to unlock large recommender model’s full potential. In the future, we aim to further investigate the efficient reasoning in larger recommender models, striving for optimal “thinking” in recommendations.
Acknowledgments and Disclosure of Funding
The work described in this paper was supported by Research Grants Council of Hong Kong(PolyU/15209724, PolyU/15205325, PolyU/15207122), and PolyU internal grants (BDWP).
Appendix A Related Work
A.1 Reinforcement Learning for LLM Reasoning
OpenAI’s GPT‑o1 demonstrated that scaling reinforcement learning (RL) training with large compute budgets enables models to internalize reasoning as a differentiable policy, achieving state-of-the-art performance in emergent meta-reasoning tasks [41]. Early RLHF methods trained a reward model on human preference data and fine-tuned the policy using Proximal Policy Optimization (PPO) [28]; however, PPO’s multiple optimization passes introduce implementation complexity. To streamline RL tuning, Direct Preference Optimization (DPO) [42] performs single-step policy updates, trading simplicity for potential off-policy bias. Alternative estimators, such as Group Relative Policy Optimization (GRPO) [27], derive baselines from group-level rewards, eliminating the need for a separate critic, while REINFORCE’s leave-one-out (RLOO) [31] computes advantages by excluding each trajectory sequentially. Building on these advancements, DeepSeek-Zero [40] removes reliance on supervised fine-tuning by driving reasoning emergence purely through intrinsic RL rewards, and DeepSeek-R1 further integrates rule-driven reward engineering with self-play verification to enhance reasoning robustness.
A.2 Large Language Model-based Recommendation
Recommendation systems are a cornerstone of modern information services [43, 44], and the integration of Large Language Models (LLMs) has emerged as a promising frontier to advance their capabilities [45, 2, 46]. Recent advances in LLM-based recommendation typically follow two paradigms: 1) LLM-enhanced, where LLMs enrich recommendation pipelines with additional features, and 2) LLM-centric, where recommendation is framed as a generative task via item identifiers [1, 2, 3, 4, 24, 25, 26], or LLMs serve as encoders to embed users and items [7, 29]. We term models under the second paradigm as large recommender models. In the first paradigm, LLM-generated embeddings are utilized to enhance conventional recommenders. Adoption of LLM-generated text are not new, typical examples includes user intent [16, 47], search queries [21], item summaries [19], or rationales [23]. Among them, SLIM [23], LangPTune [16], and Rec-R1 [21] further optimize LLM for finer rationale generation with RL. However, such designs introduce high resource demands for training and serving both models, and the connection between two modules remains gradient-less, blocking true end-to-end learning, resulting in suboptimal convergence and performance. Large recommender models typically recommend through autoregressive generation of item identifiers [4, 24] or titles [1, 2, 3, 25], or by embedding user interaction sequences [7, 29]. Large recommender model has demonstrated remarkable potential, but current methods fail to utilize reasoning ability of LLMs. To the best of our knowledge, no existing large recommender model improves recommendation over explicit reasoning generation, which is the gap that motivated our work.
Appendix B Dataset
Table 4: Dataset Statistics.
Dataset Users Items Density Interactions Train Val Test
Video Games 29,230 10,144 0.031% 63,502 50,801 6,350 6,351
CDs and Vinyl 7,701 12,024 0.023% 13,435 10,748 1,343 1,344
Musical Instruments 15,656 10,320 0.031% 34,373 27,498 3,437 3,438
@@ Appendix B Dataset
Following the temporal‐truncation protocol introduced in previous literature [3, 2], we construct three subsets—CDs and Vinyl (CDs), Video Games (Games), and Musical Instruments (Instruments) – from the lastest public Amazon review dataset [^1] https://amazon-reviews-2023.github.io/index.html. spanning from May 1996 to October 2023. For each domain, we begin with the most recent year of interactions (October 2022 – October 2023) and, if the resulting number of valid items is insufficient, iteratively roll the time window backward month-by-month until 10k items are obtained. We omit the 5-core filter so as to retain the nature behaviour characteristic of recommendation scenarios. Each user’s interaction history is then chronologically sorted and truncated to the latest 20 actions, yielding fixed-length sequences for all subsequent modelling and evaluation. Finally the dataset is split with 80% training 10% validation 10% test. The resulting statistics are listed in Table 4.
Appendix C Baselines